OpenStudy (anonymous):

Dont understand this problem: What are the solutions to x^2 = 18x. Been stuck for a while on this

OpenStudy (anonymous):

x . x = 18 x so x( x - 18 ) = 0

OpenStudy (anonymous):

the options are: A) all real numbers B) no solutions C) x = 0, x=-3 D) x=0, x=3

OpenStudy (anonymous):

which one would it be?

OpenStudy (anonymous):

E) none of the above

OpenStudy (anonymous):

theres no E... duhh

OpenStudy (anonymous):

this is why i dislike multiple choice, because people dont do the math

OpenStudy (anonymous):

i just dont know how to

OpenStudy (anonymous):

like i know x=0 but would it be C or D..where does the negatives come in?

OpenStudy (anonymous):

we start with x^2 = 18x moving everything over to 1 side x^2-18x=0 now, use the distributive property of multiplication to factor out an x x(x-18)=0 now, using the zero property of multiplication, we can conclude either x is equal to 0 or x-18 is equal to 0 giving us the 2 nice equations x=0 and x-18=0 we look at x-18=0 because its the only one thats unsolved move the 18 over to the right x=18

OpenStudy (anonymous):

i told you, the answer is not there

OpenStudy (anonymous):

so it would be no solutions?

OpenStudy (anonymous):

no.... its E) none of the above

OpenStudy (anonymous):

i solved it out, showed you the 2 solutions..... gah

OpenStudy (anonymous):

i understand..it is just that there is no E...

OpenStudy (anonymous):

then your book/teacher/website/problem/answer made a typo

OpenStudy (anonymous):

ohh..ok thanks so much..will ask about it tmrw!

OpenStudy (anonymous):

or you didnt give us the correct problem or solutions

OpenStudy (anonymous):

what does all real numbers mean?

OpenStudy (anonymous):

could that be the answer?

OpenStudy (anonymous):

"completeidiot" is right. There is an issue with the question and/or answers provided. And "all real numbers" refers to every number imaginable, positive or negative. The answer(s) to this are only x=0 and x=18. Example: What are the solution(s) to x + 3 = x + 3 ? All real numbers. Because you can plug ANY number into x and the equation will be true.

OpenStudy (anonymous):

thanks so much!